OpenStudy (anonymous):

Two groups of friends are hiking. On a coordinate grid of the area, marked in 1 mile squares, one group is at (-2, 3). Another group has hiked to (4, -1) on the same coordinate grid. If the two groups want to meet at the halfway point for lunch, to what coordinates will they travel? a. (3, 2) b. (3, -1) c. (-2, 4) d. (1, 1)

OpenStudy (anonymous):

use the midpoint formula to solve...\[\left\{\left( x _{1}+x _{2} \right)/2 , \left( y _{1}+y _{2} \right)/2 \ \right\}\]

OpenStudy (anonymous):

x=(-2+4)/2=1 y=(3-1)/2=1 (1,1)